# Kobayashi hyperbolicity of complete linear systems on abelian varieties

### v1

@@ 1. Introduction
This already improves [IMRY, Corollary 1.13], and Proposition 5.5 below gives an even better (but more involved) result. Moreover, the following question arises naturally:
Question F.
Let $(A,L)$ be a general polarized abelian variety of dimension $g\geq 2$, with
$h^{0}(A,L)>(g-1)^{g}\,.$
Is the linear system $|L|$ Kobayashi hyperbolic?
@@ 2. Algebraic hyperbolicity
They have been motivated by Fujita’s conjecture [F], and by both classical and more recent results on hyperbolicity of (very) general hypersurfaces in projective spaces (see op.cit., the references therein, and [DR]). They proved their conjecture for smooth projective toric varieties. More recently, [IMRY] produced an useful criterion to detect the algebraic hyperbolicity of basepoint-free ample complete linear systems on a variety $X$ with nef tangent bundle (and possibly mild singularities). We state it here directly when $X=A$ is an abelian variety, and refer the reader to [IMRY, Theorem 4.7] for the more general statement. Its proof builds on a nowadays standard deformation type argument, recently simplified and further developed by Coskun and Riedl in [CR1, CR2].
@@ 3. Coherent sheaves on abelian varieties: regularity and generation properties
If our line bundle $L$ has some base divisor, then we still have that (3.2) is nef, and we may take $\delta=\frac{m}{m-1}$. In this case the Criterion 2.4 applies as soon as $(m-\delta(g-2))L$ is ample, that is, as soon as
@@
one has that $|L|$ is very ample and hyperbolic. Indeed, (3.4) gives, in particular, that $L$ is globally generated, and one still has in this case, again by [JP], that (3.4) is equivalent to the $M$-regularity of $M_{L}{\langle}\frac{1}{g-2}L{\rangle}$. Then, the proof proceeds in the same way as before.
@@ 5. The basepoint-freeness threshold and hyperbolicity of linear systems
for any general polarized abelian variety $(A,L)$ of dimension $g$ and of type $(d_{1},\ldots,d_{g})$. Since $r<\frac{1}{g-1}$, we get the hyperbolicity of $|L|$ by Remark 5.1. ∎
In particular, one obtains at once Proposition E of the Introduction. Similar (and sometimes better) results can also be obtained from [Ji2, Theorem 3.1].
When $(A,L)$ is general in its moduli space, it is also expected (see [Ji2, p. 3]) that
$\beta(A,L)\sim\frac{1}{\sqrt[g]{h^{0}(A,L)}}\,.$
Therefore, the following question seems natural.
Question 5.6.
Let $(A,L)$ be a general polarized abelian variety of dimension $g\geq 2$, with
$h^{0}(A,L)>(g-1)^{g}\,.$
Is the linear system $|L|$ Kobayashi hyperbolic?

### v2

@@ 1. Introduction
This already improves [IMRY, Corollary 1.13], and Proposition 5.5 below gives an even better (but more involved) result. Moreover, the following weaker version of Conjecture D holds true when $g\leq 6$, and it has not yet been verified only for a finite number of primitive polarization types in each dimension $g\geq 7$:
Conjecture F.
Let $(A,L)$ be a general polarized abelian variety of dimension $g\geq 2$. If
$h^{0}(A,L)>\frac{(g\,(g-1))^{g}}{g!}\,$,
then $|L|$ is Kobayashi hyperbolic.
Beauville [Be] recently observed that a complete linear system $|L|$ on an abelian variety $A$ has maximal variation, provided that $L$ is an ample and globally generated line bundle on $A$. Maximal variation here means that, if $D$ is a general member in $|L|$, there are only finitely many divisors $D^{\prime}\in|L|$ which are isomorphic to $D$. Thereby, the results of the present paper yield non-trivial examples of non-isomorphic hyperbolic subvarieties of $A$.
@@ 2. Algebraic hyperbolicity
They have been motivated by Fujita’s conjecture [F], and by both classical and more recent results on hyperbolicity of (very) general hypersurfaces in projective spaces (see op.cit., the references therein, and [DR]). They proved their conjecture for smooth projective toric varieties. More recently, [IMRY] produced an useful criterion to detect the algebraic hyperbolicity of basepoint-free complete linear systems on a variety $X$ with nef tangent bundle (and possibly mild singularities). We state it here directly when $X=A$ is an abelian variety, and refer the reader to [IMRY, Theorem 4.7] for the more general statement. Its proof builds on a nowadays standard deformation type argument, recently simplified and further developed by Coskun and Riedl in [CR1, CR2].
@@ 3. Coherent sheaves on abelian varieties: regularity and generation properties
If our line bundle $L$ has some base divisor, then we still have that (3.2) is nef, and we may take $\delta=\frac{m}{m-1}$. In this case, the Criterion 2.4 applies as soon as $(m-\delta(g-2))L$ is ample, that is, as soon as
@@
one has that $|L|$ is hyperbolic. Indeed, (3.4) gives, in particular, that $L$ is globally generated, and one still has in this case, again by [JP], that (3.4) is equivalent to the $M$-regularity of $M_{L}{\langle}\frac{1}{g-2}L{\rangle}$. Then, the proof proceeds in the same way as before.
@@ 5. The basepoint-freeness threshold and hyperbolicity of linear systems
for a general polarized abelian variety $(A,L)$ of dimension $g$ and of type $(d_{1},\ldots,d_{g})$. Since $r<\frac{1}{g-1}$, we get the hyperbolicity of $|L|$ by Remark 5.1. ∎
In particular, one obtains at once Proposition E of the Introduction. Similar (and sometimes better) results can also be obtained from [Ji2, Theorem 3.1]. Moreover, the same reasoning suggests the following weaker version of Conjecture D:
Conjecture 5.6.
Let $(A,L)$ be a general polarized abelian variety of dimension $g\geq 2$. If
$h^{0}(A,L)>\frac{(g\,(g-1))^{g}}{g!}\,$,
then $|L|$ is Kobayashi hyperbolic.
Thanks to [Ji2, Theorem 1.5], Conjecture 5.6 holds true when $g\leq 6$, and there are only finitely many primitive polarization types in each dimension $g\geq 7$ for which the conjecture has not yet been verified (see [Ji2, §3] for details).
Finally, let us recall that, by [It1, Lemma 3.4(i)], for any polarized abelian variety $(A,L)$ of dimension $g$, one has that
$\frac{1}{\sqrt[g]{h^{0}(A,L)}}\leq\beta(A,L)\,$,
and, when $(A,L)$ is general in its moduli space, it is expected (see [Ji2, p. 3]) that $\beta(A,L)$ is quite close to $\frac{1}{\sqrt[g]{h^{0}(A,L)}}$. Hence, sometimes even better estimates than the one in Conjecture 5.6 may be possible.